OpenStudy (anonymous):

help solving this DE with bernoulli Equation: t^2(dy/dt) +y^2 = ty

OpenStudy (anonymous):

I got the part where I have to make it in the form to get y^n where i get dy/dt - (1/t)y = (1/t^2)y^2 but i have a hard time understanding the rest... if u = y^(1-n) then n=2 therefore u= y^(-1) and If i derive that I get du/dx = -y^-2 (dy/dx) but then I'm stuck.... what happens with the y to plug it in, I;m still stuck with it